2(3x- 7)+4(3x+ 2)= 6(5x + 9)+3
please solve and explain.

Answer by algebrahouse.com(1659)

2(3x - 7) + 4(3x + 2) = 6(5x + 9) + 3
6x - 14 + 12x + 8 = 30x + 54 + 3 {used distributive property}
18x - 6 = 30x + 57 {combined like terms}
-6 = 12x + 57 {subtracted 18x from both sides}
-63 = 12x {subtracted 57 from both sides}
x = -63/12 {divided both sides by 12}
x = -21/4 {reduced}

For more help from me, visit: www.algebrahouse.com